OpenStudy (anonymous):

lim x-sinx/x+sinx=? x->0

OpenStudy (anonymous):

plz help me!!

OpenStudy (across):

\[\lim_{x\to0}\frac{x-\sin(x)}{x+\sin(x)}=\lim_{x\to0}\frac{1-\cos(x)}{1+\cos(x)}.\]Does that make it easier? (L'Hôpital's rule)

myininaya (myininaya):

gj across

myininaya (myininaya):

lets see if we can do it without L'Hopital's way \[\lim_{x \rightarrow 0}\frac{x-\sin(x)}{x+\sin(x)} \cdot \frac{\frac{1}{x}}{\frac{1}{x}}\] \[=\lim_{x \rightarrow 0}\frac{1-\frac{\sin(x)}{x}}{1+\frac{\sin(x)}{x}}=\frac{1-1}{1+1}=0/2=0\]

OpenStudy (anonymous):

thanks a alot!